contractor plans to use four tractors to work on a project in a remote area. The
probability of a tractor functioning for a year without a break-down is known to be
80 %. If X denotes the number of tractors operating at the end of a year, determine the
probability mass and distribution functions of X.

Answer by Boreal(15235)

The mean is xp=0.8*x
the variance is xp(1-p)=0.8*0.2*x=0.16x, and the sd is 0.4*sqrt(x)